OpenStudy (anonymous):

Point A is located at (-8, 3). Point C is located at point (12, -4). Find point B if it is (2/5) the distance from point A to point B. sqrt[(12 + 8)2 + (-4 - 3)2] sqrt[(20)2 + (-7)2] sqrt[400 + 49] sqrt[449] AC = 21.18 2/5 of 21.18 is 8.472 What do I do from here?

OpenStudy (anonymous):

@phi @TuringTest

OpenStudy (phi):

it is easier to find the change in x from point A to point C: -8 to 12 is 20 units to the right. you want to go 2/5 of that amount in the x direction: 2/5 * 20= 8 add 8 to point A's x: -8+8 = 0 that is the x coordinate of point B we do the same for the y coord

OpenStudy (phi):

the y coord of A is 3 the y coord of c is -4 that is a change of -7 2/5 of -7 = -14/5 add that to 3 (3 is the same as 15/5) to get 15/5 - 14/5 = 1/5 that is the y coord of pt B

OpenStudy (anonymous):

So B is at (0, 1/5) or (0, 0.2)?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Ok, thank you! I have been stuck for like 2 hours now.